# Self-Guided Open-Vocabulary Semantic Segmentation

### 07 Dec 2023 v1

@@ 3.1 Local Region Captioning
This step is applied recursively until convergence or 8 times at most.
Caption Generation. The next step involves turning clustered, denoised and enhanced BLIP embeddings into text. The BLIP text decoder is a Transformer architecture capable of processing unordered sets of embeddings of arbitrary size. We leverage this feature and feed flattened subsets of patch embeddings, each corresponding to a cluster, to the BLIP text decoder. Spatial information is preserved due to the presence of positional embeddings added in the clustering step. With this technique, our method essentially iterates over semantic categories captured by clusters and represented by BLIP embeddings. To the best of our knowledge, we are the first to use the text decoder in this manner, enabling local captioning for which it was never specifically trained. The caption generation is stochastic, with different object namings appearing in the captions depending on random initialization. To obtain a rich, unbiased and diverse set of object names, we regenerate captions with each embedding multiple times in caption generation cycles.
@@ 3.3 Self-Guidance
The output of the clustered BLIP embeddings (see Sect. 3.1) is a 32x32 grid with cluster index assignments. Each element of the grid corresponds to an image patch from the original image, which enables us to create segmentation masks - defined as a union of areas covered by image patches with the same index - essentially for free. For example, Fig. 3 shows the clustered output as a 2-dimensional mask, which succeed in masking relevant objects to a certain extent. However, the segmentation outputs are constrained to a 32x32 resolution of the BLIP encoder. Upsampling yields unsatisfactory results, since objects appear oversegmented and boundaries remain unsharp. Another challenge is to associate each mask with one class name. In this cluster representation, masked areas are described with sentences rather than single nouns. When such a sentence contains multiple nouns, it is not trivial which noun to select as the predicted class name for that mask. To perform effective and accurate open-vocabulary semantic segmentation, we leverage Blip-Cluster-Caption’s strength in generating an elaborate set of relevant class names and use this as guidance for a pre-trained OVS model capable of producing high-resolution outputs. BLIP-Cluster-Caption is model-agnostic, therefore any kind of OVS model that expects an image and a set of class labels can be combined with our method. We focus on X-Decoder[38], which is a popular and well-performing OVS model.
@@ 3.4 Custom Open-Vocabulary Evaluator
As discussed in Sect. 2, previous works in OVS have mainly focused on a setting where the target class names are provided by the user. Hence, evaluation is possible by having access to the ground truth of those target class names during evaluation. However, in scenarios where the target class names are discovered rather than pre-specified, as in our framework, there may be a lack of direct alignment between the semantics of these categories and the classes used in the annotations. In zero-guidance segmentation [25], the authors have proposed to align class names based on the cosine-similarity in the latent space of Sentence-BERT [23] or CLIP [22]. In our initial experiments, this yielded unsatisfactory results where even in obvious cases, the open-vocabulary classes were linked to the incorrect class. Relations between two class names can be complex, such as synonymity, hyponymy, hypernymy, meronymy or holonymy, which exceed the capabilities of a cosine-similarity criterium in the respective latent spaces. As a possible solution, we propose an LLM-based Open-Vocabulary Evaluator (LOVE). Specifically, we employ the Llama 2 7B Large Language model444www.huggingface.co/meta-llama/Llama-2-7b with the task of mapping predicted open-vocabulary categories to target categories in the dataset. First, Self-SEG returns a segmentation mask with classes from an open vocabulary generated by Blip-Cluster-Caption. Second, LOVE collects all unique predicted open-vocabulary categories in the output and builds a mapper from each predicted category to the most relevant target category. Thirdly, all values in the segmentation mask belonging to each class are updated respectively according to the mapper. Finally, the mIoU is computed with the updated segmentation mask. Pseudocode and prompts for LOVE are provided in App. A.
Figure 4: Segmentation Using VLMs. Example outputs from VOC (top), ADE20K (middle), CityScapes (bottom) obtained with different configurations of our self-guidance framework, from which Self-SEG performs most accurately with automatically detected class names.
@@ 4.1 Experimental Setup
We evaluate our method on three popular semantic segmentation validation datasets: Pascal VOC [9], ADE20K [37] and CityScapes [6] with 20, 847 and 20 classes respectively. These datasets cover a wide range of difficulty and class diversity. For example, ADE20K is challenging due to a large variety of classes with many of them appearing only a few times in the whole dataset. We list dataset statistics in Tab. 3. For qualitative comparison with previous works, we use mean Intersection over Union (mIoU) as the main evaluation metric. As mentioned in Sec. 3.4, class predictions in the segmentation mask are mapped using LOVE before the mIoU is computed. To study the class-agnostic segmentation performance of different VLMs, we use class-agnostic mean Intersection over Union (cmIoU). The computation of cmIoU involves the following steps. First, pairwise IoU values, denoted as $\text{IoU}(g_{i},p_{j})$ are calculated between all ground-truth segments $g_{i}$ and predicted segments $p_{j}$ from a given image. Each $g_{i}$ segment is matched with the predicted segment $p_{\text{max}_{i}}$ that yields the highest IoU value: $p_{\text{max}_{i}}=\arg\max_{j}\text{IoU}(g_{i},p_{j})$. The cmIoU is then determined by calculating the arithmetic mean of these highest IoU values across the dataset, which can be formulated as $\text{cmIoU}=\frac{1}{N}\sum_{i=1}^{N}\text{IoU}_{\text{max}_{i}}$ where $N$ is the total number of ground-truth segments in the dataset.
In the Self-SEG model used for our experiments, we use the BLIP model built with ViT-Large backbone and finetuned for image captioning on the COCO dataset in combination with the Focal-T variant of X-Decoder. For both models, we use publicly available pretrained weights and do not perform any additional finetuning. Parameter tuning is performed for the clustering and captioning modules. Parameters of the clustering include image scales encoded by BLIP ($384\times 384$ and $512\times 512$), the $k$ values of $k$-means clustering (2 and 8), the parameters of Gaussian smoothing in the CRF (smoothness weight of 6 and smoothness $\theta$ of 0.8), the number of iterations of majority filtering (8) and the feature dimension size of the positional embeddings (256). As for text generation, we opted for nucleus sampling with a minimum length of 4 tokens and maximum of 25, top P value of 1 and repetition penalty of 100 to ensure as many unique nouns as possible. The aforementioned parameters were determined using Bayesian optimization on the Pascal VOC with the goal of maximizing the mIoU.
Figure 5: Qualitative Results. Self-SEG shows remarkable capability to identify open vocabulary class names and segment them accurately across different datasets (top, middle, bottom row are from VOC, ADE20K and CityScapes respectively).
@@ 4.2 Ablations
To investigate the performance of different configurations of our framework, we present two ablation studies. The first ablation study investigates different methods of using a VLM to obtain segmentation masks. In the second one, we examine the effect of different amounts of captioning cycles.
@@
Segmentation with Vision Language Models. A VLM can be leveraged in different ways to obtain segmentation masks. In this ablation study, we compare three methods: 1) noun-filtered captions of a plain BLIP encoder/decoder as textual guidance for X-Decoder, which then produces the final segmentation; 2) BLIP cluster embeddings from our Blip-Cluster-Caption approach used as segmentation masks (see Sect. 3.3); and 3) Self-SEG in which textual guidance is achieved through Blip-Cluster-Caption and then provided to X-Decoder. Due to the difficulty with establishing correspondence between BCC clusters and ground-truth labels, we evaluate using cmIoU. Quantitative and qualitative results are reported in Tab. 1 and Fig. 4 respectively. On all three datasets, Self-SEG performs most effectively. Comparing X-Decoder + BLIP to Self-SEG qualitatively, one can observe that the former suffers from false positives and false negatives simultaneously, which in turn serves as an explanation for the performance gap in cmIoU. For example, in Pascal VOC, it incorrectly masks the area around the bus at the cost of segmenting people. In ADE20K, it ignores large parts of the area which could appear to BLIP as insignificant. In our framework, the BLIP-Cluster-Caption ensures that smaller, less salient objects are effectively captioned. As expected, intermediate masks of BLIP-Cluster-Caption do not segment objects sufficiently accurate due to the low resolution of the embedding space. Interestingly, it is still capable of segmenting global areas linked to certain objects in Pascal VOC and ADE20K, while failing completely for CityScapes, likely due to the higher instance count.
Approach VOC [9] ADE20K [37] CityScapes [6]
X-Decoder [38] + BLIP [15] 38.0 26.7 29.2
Blip-Cluster-Caption 16.3 11.3 0.85
Self-SEG 41.9 29.2 35.8
Table 1: Ablations on Segmentation with VLMs. The cmIoU values obtained by different variants of our method. Self-SEG, which combines BCC with X-Decoder segmentation, achieves the best performance on all three datasets.
Captioning Cycles. One important feature of Self-SEG is that it can repeat the process of captioning from BLIP clusters in order to obtain a more elaborate set of nouns. This is particularly useful in situations where objects are initially missed or captioned with a suboptimal semantic description. We analyze the effect of captioning once or repeating the process 5, 10, 15, 20 or 25 times while measuring the mIoU performance. Our results, shown in Tab. 3, indicate that the optimal number of captioning cycles strongly depends on the dataset used. In particular, this number increases with the average amount of instances and unique classes per image (indicated with $\overline{I}$ and $\overline{C}$ in Tab. 3 respectively). For Pascal VOC, a dataset which contains fewer instances and unique classes per image, a single cycle is sufficient for Blip-Cluster-Caption to determine the relevant class names. It should be noted, however, that the average number of instances and classes per image are statistics obtained from the annotations of the dataset. In Pascal VOC, many objects that exist in the images are not annotated and are therefore considered background when evaluated. Hence, the actual optimal number of captioning cycles might be higher. In contrast to Pascal VOC, ADE20K, which contains on average 16 more instances and 7 unique classes per image, requires ten captioning cycles to obtain optimal performance. In the case of CityScapes, images are more complex with 34.3 instances and 17 unique classes on average. Here, twenty cycles of Blip-Cluster-Caption enable the model to obtain an mIoU score of 41.1, whereas Self-SEG would obtain 27.4 mIoU if it had used a single cycle. These outcomes indicate that increasing the number of captioning cycles with higher counts of instances and unique classes per image enable identifying more relevant objects, which in turn benefits the final segmentation output.
Method Auto-Classes VOC [9] ADE20K [37] CityScapes [6]
ZSSeg [33] 83.5 7.0 34.5
FC-CLIP [36] 95.4 14.8 56.2
X-Decoder [38] 96.2 6.4 50.8
OpenSeg [10] 72.2 8.8 -
ODISE [32] 82.7 11.0 -
OVSeg [16] 94.5 9.0 -
SAN [34] 94.6 12.4 -
CatSeg [5] 97.2 13.3 -
LSeg [14] 47.4 - -
OVSegmentor [31] 53.8 - -
SimSeg [35] 57.4 - -
OVDiff [11] 69.0 - -
ZeroSeg [25] 20.1 - -
Self-SEG (Ours) 41.6 6.4 41.1
Table 2: Quantitative Results. Self-SEG performs competitive with existing OVS methods which require class names, while achieving superior performance on the zero label setting.
4.3 Quantitative Analysis
We compare our method against existing OVS methods that require class names to be given, as well as with Zero-Guidance Segmentation (referred to as ZeroSeg [25]). Tab. 2 shows the results. Without any specification of class names, Self-SEG achieves 43%, 43% and 73% of the best performing OVS methods on Pascal VOC, ADE20K and CityScapes respectively. In the ADE20K dataset, Self-SEG is able to remarkably match the performance of X-Decoder [38] without any access to the ground truth class names, while on CityScapes it even outperforms ZSSeg [33] by 6.6 mIoU. It is noteworthy that on CityScapes, our method achieves 81% of X-Decoder’s original performance with given class names. Comparing to ZeroSeg, Self-SEG achieves superior performance on Pascal VOC (41.6 over 20.1 mIoU). Furthermore, we set a new state-of-the-art on ADE20K and CityScapes under the auto-class generation task setting. This outcome underscores the efficacy of our method in dealing with complex scenes that have an open-vocabulary nature, such as ADE20K with its large number of rare classes. While other models have to decide between 847 possibilities, Self-SEG suggests a smaller yet relevant set of class names. Our method deals well with high number of instances and can successfully identify various non-salient objects. A comparison with the findings from Tab. 3 reveals that while Blip-Cluster-Caption struggles with providing high-quality masks, it is able to identify the relevant class names, which can subsequently be segmented accurately using an OVS model. This demonstrates that the integration of Blip-Cluster-Caption and OVS harnesses the strengths of both, leading to enhanced self-guided segmentation performance.
@@ 4.3 Quantitative Analysis
Pascal VOC [9] 3.5 3.5 20 41.6 37.9 37.4 35.6 35.1 35.6
ADE20K [37] 19.5 10.5 847 5.1 5.9 6.4 5.9 6.1 6.0
CityScapes [6] 34.3 17 20 27.4 35.9 39.7 40.2 41.1 41.0
Table 3: Ablations on Captioning Cycles. The optimal number of captioning cycles strongly depends on the dataset. Datasets with more instances ($\overline{I}$) and classes on average ($\overline{C}$) benefit from multiple cycles, while captioning once for few instances suffices.

### 20 Mar 2024 v2

@@ 3.1 Local Region Captioning
This step is applied recursively until convergence or 8 times at most. In the supplementary material, we visualize the effect of each step on the embeddings.
Captioning. The next step involves turning clustered, denoised and enhanced BLIP embeddings into text. The BLIP text decoder is a Transformer architecture capable of processing unordered sets of embeddings of arbitrary size. We leverage this feature and feed flattened subsets of patch embeddings, each corresponding to a cluster, to the BLIP text decoder. Spatial information is preserved due to the presence of positional embeddings added in the clustering step. With this technique, our method essentially infers semantic categories captured by clusters and represented by BLIP embeddings. To the best of our knowledge, we are the first to use the text decoder in this manner, enabling local captioning for which it was never specifically trained. The caption generation is stochastic, with different object namings appearing in the captions depending on initialization. To obtain a rich, unbiased and diverse set of object names, we regenerate captions with each embedding multiple times in caption generation cycles.
@@ 3.3 Self-Guidance
The output of the clustered BLIP embeddings (see Sect. 3.1) is a 32x32 grid with cluster index assignments. Each element of the grid corresponds to an image patch from the original image, which enables us to create segmentation masks - defined as a union of areas covered by image patches with the same index - essentially for free. For example, Fig. 3 shows the clustered output as a 2-dimensional mask, which succeeds in masking relevant objects to a certain extent. However, the segmentation outputs are constrained to a 32x32 resolution of the BLIP encoder. Upsampling yields unsatisfactory results, since objects appear oversegmented and boundaries remain unsharp. Another challenge is to associate each mask with one class name. In this cluster representation, masked areas are described with sentences rather than single nouns. When such a sentence contains multiple nouns, it is not trivial which noun to select as the predicted class name for that mask. To perform effective and accurate auto-vocabulary semantic segmentation, we leverage BCC’s strength in generating an elaborate set of relevant class names and use this as textual guidance for a pre-trained OVS model capable of producing high-resolution outputs. BCC is model-agnostic, therefore any kind of OVS model that expects an image and a set of class labels can be combined with our method. We focus on X-Decoder[42], which is a popular and well-performing OVS model.
@@ 3.4 Custom Open-Vocabulary Evaluator
As discussed in Sect. 2, previous works in OVS have mainly focused on a setting where the target class names are provided by the user. Hence, evaluation is possible by having access to the ground truth of those target class names during evaluation. However, in scenarios where the target class names are discovered rather than pre-specified, as in our framework, there may be a lack of direct alignment between the semantics of these categories and the classes used in the annotations. In zero-guidance segmentation [28], the authors have proposed to align class names based on the cosine-similarity in the latent space of Sentence-BERT [26] or CLIP [25]. In our initial experiments, this yielded unsatisfactory results where, even in obvious cases, the open-vocabulary classes were linked to the incorrect class (e.g. taxi to road instead of car). This led to a decline in performance as measured by segmentation metrics, yet the method showed promising results from a qualitative perspective. Relations between two class names can be complex, such as synonymity, hyponymy, hypernymy, meronymy or holonymy, which exceed the capabilities of a cosine-similarity criterium in the respective latent spaces. As a possible solution, we propose an LLM-based Auto-Vocabulary Evaluator (LAVE). Specifically, we employ the Llama 2 7B Large Language model444www.huggingface.co/meta-llama/Llama-2-7b with the task of mapping predicted auto-vocabulary categories to target categories in the dataset. First, AutoSeg returns a segmentation mask with classes from an auto-vocabulary generated by BCC. Second, LAVE collects all uniquely predicted auto-vocabulary categories in the output and builds a mapper from each predicted category to the most relevant or similar target category. Thirdly, all pixel values in the segmentation mask belonging to each class are updated respectively according to the mapper. Finally, the mean Intersection over Union (mIoU) is computed with the updated segmentation mask. Pseudocode and prompts for LAVE are provided in the supplementary material.
Figure 4: Segmentation with VLMs. Example outputs on PASCAL VOC/Context (top), ADE20K (middle) and Cityscapes (bottom) by feeding plain BLIP embeddings to X-Decoder, directly using BCC embeddings as masks and AutoSeg. Notably, our method segments images in the most comprehensive and accurate manner.
@@ 4.1 Experimental Setup
We evaluate our method on three popular semantic segmentation validation datasets: PASCAL VOC [9] and Context [23], ADE20K [41] and Cityscapes [6] with 20, 459, 847 and 20 classes respectively. These datasets cover a wide range of difficulty and class diversity. For example, ADE20K is challenging with its many and infrequently appearing classes, while Cityscapes contains many object instances per image. We list dataset statistics in Tab. 2.
Evaluation. For quantitative comparison with previous works, we use mIoU as the main metric. As mentioned in Sec. 3.4, class predictions in the segmentation mask are mapped using LAVE before the mIoU is computed. Furthermore, to study the class-agnostic segmentation performance of different VLMs, we report class-agnostic mean Intersection over Union (cmIoU). The computation of cmIoU involves the following steps. First, pairwise IoU values, denoted as $\text{IoU}(g_{i},p_{j})$ are calculated between all ground-truth segments $g_{i}$ and predicted segments $p_{j}$ from a given image. Each $g_{i}$ segment is matched with the predicted segment $p_{\text{max}_{i}}$ that yields the highest IoU value: $p_{\text{max}_{i}}=\arg\max_{j}\text{IoU}(g_{i},p_{j})$. The cmIoU is then determined by calculating the arithmetic mean of these highest IoU values across the dataset, which can be formulated as $\text{cmIoU}=\frac{1}{N}\sum_{i=1}^{N}\text{IoU}_{\text{max}_{i}}$ where $N$ is the total number of ground-truth segments in the dataset.
Table 1: Ablations on Segmentation with VLMs. Class-agnostic mIoU values for different variants of our method. AutoSeg, which combines BCC with X-Decoder segmentation, achieves the best performance on all four datasets.
Approach VOC [9] PC-459 [23] ADE20K [41] Cityscapes [6]
X-Decoder [42] + BLIP [16] 38.0 35.3 26.7 29.2
Blip-Cluster-Caption 16.3 16.3 11.3 0.85
AutoSeg 71.8 47.7 29.2 35.8
Implementation details. For our experiments, we use the BLIP model built with ViT-Large backbone and finetuned for image captioning on the COCO dataset in combination with the Focal-L variant of X-Decoder. For both models, we use publicly available pretrained weights and do not perform any additional finetuning. Parameter tuning is performed for the clustering and captioning modules. Parameters of the clustering include image scales encoded by BLIP ($384\times 384$ and $512\times 512$), the $k$ values of $k$-means clustering (2 to 8), the parameters of Gaussian smoothing in the CRF (smoothness weight of 6 and smoothness $\theta$ of 0.8), the number of iterations of majority filtering (8) and the feature dimension size of the positional embeddings (256). As for text generation, we opted for nucleus sampling with a minimum length of 4 tokens and maximum of 25, top P value of 1 and repetition penalty of 100 to ensure as many unique nouns as possible. These parameters were determined using Bayesian optimization on VOC with the goal of maximizing the mIoU.
@@ 4.2 Ablations
To investigate the performance of different configurations of our framework, we present two ablation studies. The first ablation study investigates different methods of using a VLM to obtain segmentation masks. In the second, we examine the effect of different amounts of captioning cycles.
@@
Segmentation with Vision Language Models. A VLM can be leveraged in different ways to obtain segmentation masks. In this ablation study, we study the benefit of using an auto-vocabulary generated with BCC for segmentation as in AutoSeg. In order to do so, we compare our method against a scenario where plain BLIP embeddings are used for captioning instead of enhanced ones. Furthermore, we also evaluate the extent to which BCC embedding outputs themselves can be used directly as segmentation masks (see Sect. 3.3) after upsampling. Due to the difficulty of establishing correspondence between BCC clusters and ground-truth labels, we evaluate the segmentation quality using cmIoU. Quantitative results are reported in Tab. 1. On all four datasets, AutoSeg performs most effectively. Comparing X-Decoder + BLIP to AutoSeg qualitatively, one can observe that the former suffers from false positives and false negatives simultaneously, which in turn serves as an explanation for the performance gap in cmIoU. For example, in PASCAL, it incorrectly masks the area around the bus at the cost of segmenting people. In ADE20K and Cityscapes, it ignores large parts of the area which could appear to BLIP as insignificant. Our framework ensures that smaller, less salient objects are effectively captioned. As expected, intermediate masks of BCC do not segment objects sufficiently accurate due to the low resolution of the embedding space. Interestingly, it is still capable of segmenting global areas linked to certain objects in PASCAL and ADE20K, while failing completely for Cityscapes, likely due to the higher instance count and higher resolution of images.
Table 2: Ablations on Captioning Cycles. The optimal number of captioning cycles varies per dataset. On Cityscapes, which has images in the highest resolution, most instances ($\overline{I}$) and classes on average ($\overline{C}$), AutoSeg benefits from 25 cycles, while between 1 and 15 captioning cycles for datasets with fewer instances per image suffices.
@@
PASCAL VOC [9] 3.5 3.5 20 85.3 84.4 87.1 85.7 83.3 83.3
PASCAL Context [23] 18.9 6.2 459 10.1 9.7 11.2 11.7 9.6 10.8
ADE20K [41] 19.5 10.5 847 5.9 5.8 5.1 5.2 5.1 5.6
Cityscapes [6] 34.3 17 20 28.2 28.1 28.4 28.3 27.9 29.5
Captioning Cycles. A key aspect of our model is its capability to iteratively repeat the captioning process within BCC, enabling it to generate a more exhaustive collection of nouns which accurately describe the scene. This approach proves especially beneficial in instances where objects are initially overlooked or described with less precise semantic terms. To assess the impact, we explore the effects of single versus multiple captioning iterations — 5, 10, 15, 20, or 25 times — using the mIoU metric. The findings, presented in Tab. 2, reveal that the optimal number of captioning cycles varies significantly with the dataset in question. Specifically, we observe that for the Cityscapes dataset, characterized by its high average number of instances and unique classes per image (denoted as $\overline{I}$ and $\overline{C}$ in Tab. 2, respectively), a greater number of cycles yields the most benefit. For datasets with a lower average number of instances and unique classes per image, like VOC or ADE20K, fewer captioning cycles are required for BCC to identify all relevant class names. As anticipated, our model requires a greater number of captioning cycles for the Context dataset, which, despite featuring images identical to those in VOC, includes annotations for an additional 439 object categories. Remarkably, AutoSeg successfully handles ADE20K, the dataset with the largest number of unique classes in our experiments, using merely a single captioning cycle. This outcome underscores our model’s ability in efficiently addressing complex datasets which resemble an open-ended setting.
Table 3: Quantitative Results (mIoU). AutoSeg performs competitive with existing OVS methods which require class names, while achieving superior performance on the zero label setting. Results with dashes are not published and unavailable.
Method Auto-Classes VOC [9] PC-459 [23] ADE20K [41] Cityscapes [6]
LSeg [15] 47.4 - - -
OVSegmentor [36] 53.8 20.4 - -
SimSeg [39] 57.4 26.2 - -
OVDiff [11] 69.0 31.4 - -
OpenSeg [10] 72.2 45.9 8.8 -
OVSeg [17] 94.5 55.7 9.0 -
ODISE [34] 82.7 57.3 11.0 -
SAN [37] 94.6 57.7 12.4 -
CAT-Seg [5] 97.2 62.0 13.3 -
ZSSeg [38] 83.5 47.7 7.0 34.5
X-Decoder [42] 96.2 65.1 6.4 50.8
FC-CLIP [40] 95.4 58.4 14.8 56.2
ZeroSeg [28] 20.1 11.4 - -
AutoSeg (Ours) 87.1 11.7 5.9 29.5
4.3 Quantitative Analysis
We compare our method against existing OVS methods that require class names to be given, as well as with Zero-Guidance Segmentation (referred to as ZeroSeg [28]). Tab. 3 shows the results. Without any specification of class names through user input, AutoSeg achieves 86%, 18%, 40% and 52% of the best performing OVS methods on VOC, Context, ADE20K and Cityscapes respectively. The difference in performance can be attributed to AutoSeg needing to identify the subjects for segmentation independently, while OVS methods are explicitly instructed on the specific objects to segment. Remarkably, however, AutoSeg is able to remarkably surpass the performance many OVS methods on VOC, such as OVDiff [11], OpenSeg [10] or ODISE [34], without any access to the ground truth class names. While still leaving room for improvement, our model compares competitively with OVS methods on ADE20K and Cityscapes, two very challenging datasets either high in number of unique classes or instances. Comparing to ZeroSeg, AutoSeg achieves superior performance on VOC (87.1 over 20.1 mIoU) and Context (11.7 over 11.4 mIoU). Furthermore, we set a new state-of-the-art on ADE20K and Cityscapes under the auto-class generation task setting. This outcome underscores the efficacy of our method in dealing with complex scenes that have an open-ended nature, such as ADE20K with its large number of rare classes. While other models have to decide between 847 possibilities, AutoSeg suggests a smaller yet more relevant set of class names. Our method deals well with high numbers of instances and can successfully identify various non-salient objects. A comparison with the findings from Tab. 2 reveals that while BCC struggles with providing high-quality masks on its own, it is able to identify the accurate target classes, which can subsequently be segmented with precision. Additional experiments are provided in the supplementary material, in which we generate class names with other image captioning methods and compare the segmentation performance of such vocabularies. Overall, our results demonstrate that the integration of BCC and OVS harnesses the strengths of both, leading to enhanced open-ended recognition capabilities.
Figure 5: Qualitative Results. AutoSeg shows remarkable capability to identify out-of-vocabulary categories, such as hawk or coke, and segment them accurately across different datasets. The first two rows highlight results on from PASCAL VOC/Context, while the bottom two rows are from ADE20K and Cityscapes respectively.